# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v1

@@ Abstract
One of the main challenges of N-Version Programming is development cost: it requires paying multiple teams to develop variants of the same system. To address this issue, we propose the automated generation of variants using large language models. We design, develop and evaluate Galápagos: a tool for generating program variants using LLMs, validating their correctness and equivalence, and using them to assemble N-Version binaries. We evaluate Galápagos by creating N-Version components of real-world C code. Our original results show that Galápagos can produce program variants that are proven to be functionally equivalent, even when the variants are written in a different programming language. Our systematic diversity measurement indicate that functionally equivalent variants produced by Galápagos, are statically different after compilation, and present diverging internal behavior at runtime. We demonstrate that the variants produced by Galápagos can protect C code against real miscompilation bugs which affect the Clang compiler. Overall, our paper shows that producing N-Version software can be drastically automated by advanced usage of practical formal verification and generative language models.
@@ I Introduction
The largest challenge of N-Version programming relates to the development and operation costs of the approach [3]. There are high costs induced by the development and verification of multiple program versions instead of a single one, typically linear in the number of versions. Another important additional cost is the development of additional software to orchestrate the multiple versions (the N-version harness).
Large Language Models (LLMs) are powerful tools able to carry out complex source code-related tasks, such as code completion [4], automatic program repair [5], and code summarization [6]. Our key intuition is that they are able to mitigate the major cost problem of N-Version programming. Our idea is that, given a reference program, we can leverage the creative and search power of LLMs to automatically produce program variants. We anticipate these variants to be diverse, due to the great volume of data used in LLM training, apt to being combined in an N-Version system. In other words, we want to harness the LLM’s ability to synthesize program variants, and instead of having multiple teams producing multiple versions at a high costs, having the LLM producing them at virtually no cost.
In this paper, we design, implement and evaluate Galápagos, a tool for automated, verified N-Version programming. Galápagos is made up of three passes and uses off-the-shelf LLMs and leverages program equivalence verification to provide strong guarantees. Galápagos generates variants of a reference function and then assembles N-Version implementations of that function. We implement Galápagos in the scope of C libraries, with support for function variant generation as well as N-Version in same-language (C-to-C), and cross-language (C-to-Go).
@@
We conduct a systematic evaluation of Galápagos to quantify the correctness, diversity, and usefulness of both the function variants and the N-Version components. First, we assemble a dataset of 30 real-world C functions, extracted from 6 notable open source projects related to cryptography and multimedia. We run Galápagos for transforming each function in the dataset into an N-Version component. Second, we measure the diversity of the variants, both statically and dynamically. Statically, we measure the uniqueness of the resulting machine code after compiling the variants, using different optimization configurations. Dynamically, we execute each unique variant, and record the CPU instructions they utilize. This allows us to quantify the variation in execution paths, which is known to be important against certain types of exploits such as side-channel attacks [7]. Third, we rigorously test the assembled N-Version implementations in the context of miscompilation bugs [8]. It clearly demonstrates Galápagos’s ability to enhance robustness wrt. our target fault model.
@@ III-D Validation Pass
Equivalence check: In this final check, Galápagos calls a configurable off-the-shelf formal equivalence checking tool, such as alive2 [30], or Rust’s Kani [31]. This uncouples Galápagos from any specific IR language and allows the use of any state-of-the-art solution for verification. It performs the equivalence check with the reference and variant IR codes as parameters. The variant is filtered out if the tool proves the variant to be non-equal, or if it fails to prove equivalence within a given time and memory budget. Reasons for failure at this step are direct evidence of non-equivalence, as the variant deviates from the specification’s functionality. This is the most resource-consuming check, so it is executed last. All function variants that pass the equivalence check are considered correct and can be forwarded to the next pass for assembly.
@@ III-E Harnessing Pass
The previous passes of Galápagos aim at maximizing diversity. They leverage the diversity of source languages and blend multiple toolchains. Therefore, the variants’ IR contain subtle differences, which is at the core of the software diversity we are seeking. Yet, these differences need to be dealt with when assembling N-version functions. To handle this case, Galápagos performs a non-trivial IR conversion process.
Specifically, Galápagos matches function signatures and removes language-specific components. For instance, if the original project is written in C and the variant is written in Go, a set of predefined rules is applied where: (1) the function’s nest self-reference is removed; (2) the function’s parameter names are normalized; (3) IR calls to the Go "lifetime" modules are removed, and; (4) IR calls to the Go "panic" modules are replaced with crashes specific to the IR.
@@ IV-E Methodology for RQ3: Validating the Ability to Mitigate Miscompilations
As described in subsection III-A, our work specifically focuses on hardening against miscompilation errors produced by bugs in compilers, which is a very dangerous and tricky class of bugs. In RQ3, we consider miscompilation bugs from the Clang compiler. We assess the effectiveness of Galápagos at automatically generating N-Version programs which can mitigate these bugs. We look at all bugs reported in Table III, coming from different Clang versions and optimization stages. They are all unresolved in the repository’s issue tracker at the time of experimentation. For each bug, we create two test programs: one with a function containing a minimal reproducible example for said bug, and one with an N-Version implementation of the same function. We always run Galápagos with the minimal function as the reference broken run. We consider Galápagos to be successful if the N-Version implementation of the reference forces a crash instead of producing the incorrect output due to the miscompilation bug.
@@ V-A Validating Variants
Project Function Name Same-language (C) Cross-language (Go)
@@
icbrt64 10 10 10 10 9 9 9 9
BitDeinterleave 10 10 10 3 10 10 10 3
BitInterleave 10 10 10 4 1 1 1 1
@@
The average rate of success decreases for test validation and equivalence verification to 82,6% (248/300) and 40,6% (122/300), respectively. Testing fails in 40 out of 288 remaining cases, either because of at least one failed test case or test suite timeout. An example from ffmpeg’s flac_get_max_frame_size is shown in 7, where constant addition to the count is not performed correctly. Finally, equivalence validation fails in 126 out of 248 remaining cases. Here, we uncover the cases where the LLM is able to synthesize variants that successfully run and behave similarly to the original, yet the complete semantics of the reference function are not correctly captured. Figure 8 shows an equivalence validation failure example for one of the variants of function fBlaMka. The figure shows a snippet of alive-tv’s that indicates that for inputs x and y of 131 074 and 51 111 936 respectively, the reference function outputs 13 398 943 041 538, while the variant outputs 7 235 012 610.
@@
When executing the variants, the average rate of success drops for test validation and equivalence verification to 69,0% (207) and 37,3% (112/300), respectively. Testing fails in 22 out of 229 remaining cases, either because of at least one failed test case or test suite timeout. The acquired errors are logical ones, such as the one displayed in Figure 7, with no major LLM hallucinations detected in the variants generated in our experiments. This is, no nonsensical, or completely off-the-mark solutions. Finally, equivalence validation fails in 95 out of 207 remaining cases, where the LLM misses a part of the semantics of the reference function.
It is worth noting that the suitability of variants is similar in both the same-language and cross-language configurations, especially when contrasting the final number of variants that are proved to be equivalent. Furthermore, both configurations seem to display similar behavior regarding the generation of equivalent variants: both same- and cross-language seem to excel and face difficulties within the same functions. libsodium’s conversion functions are a clear example of this, consisting of a single line with several chained operations that appear to confuse the LLM, regardless of configuration.
@@
[style=mpdframe, frametitle=Answer to RQ1] Large language models successfully generate variants given a reference function. Out of 600 function variants, 234 were verifiably equivalent, 122 of which were created from same-language source code and 112 from a source code in a different language. This signals compelling applicability of creating automated cross-language N-Version functions with strong formal guarantees.
@@ V-B Static Uniqueness
For the same-language configuration, 115 of the 122 equivalent variants are different from each other at the IR level. These 115 variants are generated from 21 out of the 30 reference functions. The compilation with "-O0" preserves 115 unique variants. Meanwhile, the number of unique variants is reduced to 90 after "-O3" optimization, representing 73.77% of all equivalent variants. The optimizations observed in the functions are concerned with inlining, loop invariant code motion, and global value numbering.
For the cross-language configuration, a total of 51 unique variants were found at the IR level, among the 120 equivalent variants.
These 51 variants correspond to 19 out of 30 reference functions. On compilation, the number of unique variants is reduced to 36 after "-O3" optimization, representing 32.14% of all equivalent variants.
A key result for RQ2 is the fact that the introduced diversity is generally well-preserved across optimization stages: from all 166 unique variants at the IR level, a total of 126 (75.9%) are still unique after applying all optimization passes.
In total, at least 1 unique, equivalent variant was found for 24 out of 30 reference functions. Furthermore, out of all the functions for which we successfully synthesize at least 1 equivalent variant, we are able to find at least one unique variant. In the case of same-language diversification, the number of unique variants per function is greater than 2 in most cases. Cross-language diversification is equally good at generating unique variants, although the total number of variants is lower. We hypothesize that the lower number of unique variants in the cross-language configuration is caused by two reasons: The LLM used in our experiment produces less diverse source code given our experimental parameters, or; the Go to IR compiler produces more consistent code across variants.
Project Function Name Same-language (C) Cross-language (Go) Totals
@@
alaw_to_s16 1 1 1 1 1 1 6 4 3 3 3 3 5 4
iec958_parity 8 7 7 7 6 6 10 7 5 4 4 4 14 10
add 9 9 9 6 6 6 1 1 1 1 1 1 10 7
ulaw_to_s16 9 9 8 1 1 1 10 2 1 1 1 1 11 2
@@
ctz 10 10 10 10 10 10 10 9 8 8 8 8 19 18
int16_t_negative_mask 7 5 5 5 5 5 5 3 2 2 2 2 8 7
int16_t_nonzero_mask 7 7 6 6 6 6 5 4 1 1 1 1 11 7
@@
liboqs int16_nonzero_mask 3 3 2 2 2 2 8 3 1 1 1 1 6 3
libsodium fBlaMka 9 9 5 5 5 5 9 5
icbrt64 10 8 5 5 5 5 9 5 4 4 4 4 13 9
BitDeinterleave 3 3 3 3 3 3 3 1 1 1 1 1 4 4
BitInterleave 4 4 4 4 4 4 1 1 1 1 1 1 5 5
@@ V-C Dynamic Uniqueness
Now, we execute the diverse, equivalent variants generated by Galápagos. On average, same-language variants’ instruction sets have a Jaccard similarity coefficient of 0.788 with respect to the reference functions’ instruction sets; while the same metric for cross-language variants’ instruction sets is 0.418. In this measurement, the Jaccard similarity coefficient indicates overlap in observable internal behavior. Hence, cross-language variants are strictly better at maximizing diversity at runtime.
Figure 10 shows plots with CPU instruction counts of 6 of the dataset’s functions, one function per project. Each sub-figure shows 3 overlapping areas, aggregating the instruction counts of the original function, an equivalent same-language variant, and an equivalent cross-language variant. The exception is function fBlaMka from the libsodium project, which only shows the original function, and an equivalent same-language variant, given that no equivalent cross-language variant was found. In detail, each figure is a radial plot, where the axes represent each of the CPU instructions executed by the function variants. Each colored area aggregates the instruction count of a different function implementation. This means that the less overlap in the figures, the more diverse the variants are in terms of the number and type of executed CPU instructions.
On these plots, we observe that the cross-language variants consistently show more diverse behavior in terms of CPU instructions. For instance, the sub-figure of function icbrt64 shows an almost complete overlap of the original implementation and the same-language variant’s areas, while the cross-language variant area clearly displays a different set of instructions and count of the overlapping instructions. Similarly, in the case of function mix, the cross-language variant executes 6 instructions that are executed neither in the reference function nor in the same-language variant, while also avoiding 2 instructions from the original execution set. This difference in instruction subsets follows naturally from using different compiler front-ends for producing the variants’ IR code. Only in a few instances, the set of instructions of same-language variants differs from the original. For example, in the sub-figure corresponding to function int16_non_zero_mask, only 4 out of 12 instructions overlap, between the original and the C variant.

### v2

@@ Abstract
N-Version Programming is a well-known methodology for developing fault-tolerant systems. It achieves fault detection and correction at runtime by adding diverse redundancy into programs, minimizing fault mode overlap between redundant program variants. In this work, we propose the automated generation of program variants using large language models. We design, develop and evaluate Galápagos: a tool for generating program variants using LLMs, validating their correctness and equivalence, and using them to assemble N-Version binaries. We evaluate Galápagos by creating N-Version components of real-world C code. Our original results show that Galápagos can produce program variants that are proven to be functionally equivalent, even when the variants are written in a different programming language. Our systematic diversity measurement indicates that functionally equivalent variants produced by Galápagos, are statically different after compilation, and present diverging internal behavior at runtime. We demonstrate that the variants produced by Galápagos can protect C code against real miscompilation bugs which affect the Clang compiler. Overall, our paper shows that producing N-Version software can be drastically automated by advanced usage of practical formal verification and generative language models.
@@ I Introduction
The largest challenge of N-Version programming relates to the development and operation costs of the approach [3]. There are high costs induced by the manual development and verification of multiple program versions instead of a single one, typically linear in the number of versions. Another important additional cost is the development of additional software to orchestrate the multiple versions (the N-version harness).
Large Language Models (LLMs) are powerful tools able to carry out complex source code-related tasks, such as code completion [4], automatic program repair [5], and code summarization [6]. Our key intuition is that they are able to efficiently automate N-Version programming. Given a reference program, we can leverage the creative and search power of LLMs to automatically produce program variants. We anticipate these variants to be diverse, due to the great volume of data used in LLM training, apt to being combined in an N-Version system. In other words, we want to harness the LLM’s ability to synthesize program variants, and instead of having multiple teams producing multiple versions at a high costs, having the LLM producing them at virtually no cost.
In this paper, we design, implement and evaluate Galápagos, a tool for automated, verified N-Version programming. Galápagos is made up of three passes, which use off-the-shelf LLMs and harness program equivalence verification to provide strong guarantees. Galápagos generates variants of a reference function and then assembles N-Version implementations of that function. We implement Galápagos in the scope of C libraries, with support for function variant generation as well as N-Version in same-language (C-to-C), and cross-language (C-to-Go).
@@
We conduct a systematic evaluation of Galápagos to quantify the correctness, diversity, and usefulness of both the function variants and the N-Version components. First, we assemble a dataset of 30 real-world C functions, extracted from 6 notable cryptography and multimedia open source projects. We run Galápagos to transform each function in the dataset into an N-Version component. Second, we measure the diversity of the variants, both statically and dynamically. Statically, we measure the uniqueness of the resulting machine code after compiling the variants, using different optimization configurations. Dynamically, we execute each unique variant, and record the CPU instructions they utilize. This allows us to quantify the variation in execution paths, which is known to be important against certain types of exploits such as side-channel attacks [7]. Third, we rigorously test the assembled N-Version implementations in the context of miscompilation bugs [8]. It clearly demonstrates Galápagos’s ability to enhance robustness wrt. our target fault model.
@@ III-D Validation Pass
Equivalence check: In this final check, Galápagos calls a configurable off-the-shelf formal equivalence checking tool, such as alive2 [30], or Rust’s Kani [31]. This uncouples Galápagos from any specific IR language and allows the use of any state-of-the-art solution for verification. It performs the equivalence check with the reference and variant IR codes as parameters. The variant is filtered out if the tool proves the variant to be non-equal, or if it fails to prove equivalence within a given time and memory budget. Reasons for failure at this step are direct evidence of non-equivalence, as the variant deviates from the specification’s functionality. This is the most resource-consuming check, so it is executed last. All function variants that pass the equivalence check are considered correct and can be forwarded to the next pass for assembly. Addressing the accuracy of the equivalence checking tool is out of the scope of this work. We point that equivalence check false negatives would not introduce errors to the system, as the variant would be filtered out. Equivalence check false positives are more challenging, as non-equivalent variants would be accepted in the N-version assembly, Yet, their divergent output would not propagate silent corruptions, but instead would trigger a program termination (see  subsection III-E).
@@ III-E Harnessing Pass
The previous passes of Galápagos aim at maximizing diversity. They leverage the diversity of source languages and blend multiple toolchains. Therefore, the variants’ IR contain subtle differences, which is at the core of the software diversity we are seeking. Yet, these differences need to be dealt with when assembling N-version functions. To handle this case, Galápagos supports non-trivial IR transformations of the variants during the harnessing pass. Specifically, Galápagos implements Go-to-C transformations related to function signatures and replacing language-specific components: (1) the LLVM function’s nest self-reference is removed; (2) the function’s parameter names are normalized; and (3) IR calls to the Go "runtime" module are removed  [^1] https://github.com/ASSERT-KTH/Galapagos/blob/main/linker/lib/go/go.cpp.
@@ 3. RQ3
  4. RQ4
    What is the performance overhead introduced by the N-Version functions generated by Galápagos?
    N-Version functions with high execution overhead might be impractical. Therefore, we measure the resulting performance overhead introduced by Galápagos. To do this, we run the original and N-Version functions and compare their average execution times.
@@ IV-E Methodology for RQ3: Validating the Ability to Mitigate Miscompilations
As described in subsection III-A, our work specifically focuses on hardening against miscompilation errors produced by bugs in compilers, which is a very dangerous and tricky class of bugs. In RQ3, we consider miscompilation bugs from the Clang compiler. We assess the effectiveness of Galápagos in automatically generating N-Version programs which can mitigate these bugs. We look at all bugs reported in Table III, coming from different versions of Clang and optimization stages. All are unresolved in the repository’s issue tracker at the time of experimentation. For each bug, we create two test programs: one with a function containing a minimal reproducible example for said bug, and one with an N-Version implementation of the same function. We always run Galápagos with the minimal function as the reference broken run. We consider Galápagos to be successful if the N-Version implementation of the reference forces a crash instead of producing the incorrect output due to the miscompilation bug.
IV-F Methodology for RQ4: Measuring Performance Overhead
To measure the performance overhead of N-Version functions generated by Galápagos, we use the clock_gettime(3) API. To get an accurate comparison, we time both the original single-version and the N-Version functions over 2 000 000 executions. Execution and measurements are done in an Intel i7-11800H system, with 16 GB of RAM.
@@ V-A Validating Variants
Project Function Name Same-language diversification (C) Cross-language diversification (C to Go)
@@
icbrt64 10 10 10 7 9 9 9 0
BitDeinterleave 10 10 10 0 10 10 10 3
BitInterleave 10 10 10 0 1 1 1 1
@@
The average rate of success decreases for test validation and equivalence verification to 82,6% (248/300) and 32,3% (93/300), respectively. Testing fails in 40 out of 288 remaining cases, either because of at least one failed test case or test suite timeout. An example from ffmpeg’s flac_get_max_frame_size is shown in 7, where constant addition to the count is not performed correctly. Finally, equivalence validation fails in 151 out of 248 remaining cases. Here, we uncover the cases where the LLM is able to synthesize variants that successfully run and behave similarly to the original, yet the complete semantics of the reference function are not correctly captured. Figure 8 shows an equivalence validation failure example for one of the variants of function fBlaMka. The figure shows a snippet of alive-tv’s that indicates that for inputs x and y of 131 074 and 51 111 936 respectively, the reference function outputs 13 398 943 041 538, while the variant outputs 7 235 012 610.
@@
When executing the variants, the average rate of success drops for test validation and equivalence verification to 69,0% (207/300) and 24,3% (73/300), respectively. Testing fails in 22 out of 229 remaining cases, either because of at least one failed test case or test suite timeout. The acquired errors are logical ones, such as the one displayed in Figure 7, with no major LLM hallucinations detected in the variants generated in our experiments. This is, no nonsensical, or completely off-the-mark solutions. Finally, equivalence validation fails in 110 out of 207 remaining cases, where the LLM misses a part of the semantics of the reference function.
It is worth noting that the acceptability of variants is close for both same-language and cross-language configurations, especially when contrasting the final number of variants that are proved to be equivalent.
Furthermore, for most functions (21/30), both configurations generate a similar amount of equivalent variants. Both same- and cross-language seem to excel or face difficulties for these functions: libsodium’s conversion functions are a clear example of this, consisting of a single line with several chained operations; they appear to confuse the LLM, regardless of the diversification configuration. The remaining functions (9/30) show disparate diversification results across configurations: while for some functions the same-language configuration performs better (5/9), for others the cross-language configuration produces more equivalent variants (4/9).
We now analyze the distribution of equivalent variants per function. In the same-language configuration, for the bottom quartile, 0 equivalent variant were found; for the median, 2 equivalent variants were found; and the top quartile, 7 equivalent variants were found. In the cross-language configuration, for the bottom quartile 0 equivalent variant were found; for the median 1, equivalent variant was found; and for the top quartile, 5 equivalent variants. We can observe that both distributions are skewed towards zero, this means that non-equivalent variants are more likely to be generated.
However, it is notable that for most (23/30) of the functions, at least 1 equivalent variant was generated. We consider this to be a promising result that shows the feasibility of the Galápagos concept. We believe this can be further improved by using more advanced models and sophisticated prompting techniques.
@@
[style=mpdframe, frametitle=Answer to RQ1] Large language models successfully generate variants given a reference function. Out of 600 function variants, 170 were verifiably equivalent, 97 of which were created from same-language source code and 73 from a source code in a different language. This signals compelling applicability of creating automated cross-language N-Version functions with strong formal guarantees.
@@ V-B Static Uniqueness
For the same-language configuration, 91 of the 93 equivalent variants are different from each other at the IR level. These 91 variants are generated from 18 out of the 30 reference functions. The compilation with "-O0" preserves 80 unique variants. Meanwhile, the number of unique variants is reduced to 72 after "-O3" optimization, representing 74.22% of all equivalent variants. The optimizations observed in the functions are performed at the "-O1" level, and are concerned with inlining, loop invariant code motion, and global value numbering.
For the cross-language configuration, a total of 36 unique variants were found at the IR level, among the 73 equivalent variants.
These 31 variants correspond to 16 out of 30 reference functions. On compilation, the number of unique variants is reduced to 21 after "-O3" optimization, representing 28.76% of all equivalent variants.
A key result for RQ2 is the fact that the introduced diversity is generally well-preserved across optimization stages: from all 127 unique variants at the IR level, a total of 93 (73.22%) are still unique after applying all optimization passes.
In total, at least 1 unique, equivalent variant was found for 24 out of 30 reference functions. Furthermore, we are able to find at least one behaviorally unique variant for each function for which we successfully synthesize at least 1 equivalent variant, . In the case of same-language diversification, the number of unique variants per function is greater than 2 in most cases (15/19). Cross-language diversification, on the other hand, does not perform as well, with only one unique variant for most functions (13/16). We hypothesize that the lower number of unique variants in the cross-language configuration is caused by two reasons: The LLM used in our experiment produces less diverse source code given our experimental parameters, or; the Go-to-IR compiler produces more consistent code across variants.
Project Function Name Same-language diversification (C) Cross-language diversification (C to Go) Totals
@@
alaw_to_s16 1 1 1 1 1 1 6 4 3 2 2 2 5 3
iec958_parity 4 3 3 3 3 3 4 4 3 2 2 2 7 5
add 9 9 9 9 9 9 1 1 1 1 1 1 10 10
ulaw_to_s16 9 9 8 8 8 8 10 2 1 1 1 1 11 9
@@
ctz 9 9 9 9 9 9 6 5 4 4 4 4 14 13
int16_t_negative_mask 7 5 5 1 1 1 5 3 2 1 1 1 8 2
int16_t_nonzero_mask 7 7 6 4 4 4 5 4 1 1 1 1 11 5
@@
liboqs int16_nonzero_mask 3 3 2 1 1 1 8 3 1 1 1 1 6 2
libsodium fBlaMka 9 9 6 6 6 6 9 6
icbrt64 7 5 3 2 2 2 5 2
BitDeinterleave 3 1 1 1 1 1 1 1
BitInterleave 1 1 1 1 1 1 1 1
@@ V-C Dynamic Uniqueness
Now, we execute the diverse, equivalent variants generated by Galápagos. On average, same-language variants’ instruction sets have a Jaccard similarity coefficient of 0.805 with respect to the reference functions’ instruction sets; while the same metric for cross-language variants’ instruction sets is 0.693. In this measurement, the Jaccard similarity coefficient indicates overlap in observable internal behavior. Hence, cross-language variants are strictly better at maximizing diversity at runtime.
Figure 10 shows plots with CPU instruction counts of 6 of the dataset’s functions, one function per project. Each sub-figure shows 3 overlapping areas, aggregating the instruction counts of the original function, an equivalent same-language variant, and an equivalent cross-language variant. In detail, each figure is a radial plot, where the axes represent each of the CPU instructions executed by the function variants. Each colored area aggregates the instruction count of a different function implementation. This means that the less overlap in the figures, the more diverse the variants are in terms of the number and type of executed CPU instructions.
On these plots, we observe that the cross-language variants consistently show more diverse behavior in terms of CPU instructions. For instance, the sub-figure of function _booth_recode_w5 shows an almost complete overlap of the original implementation and the same-language variant’s areas, while the cross-language variant area clearly displays a different set of instructions and count of the overlapping instructions. Similarly, in the case of function ctz, the cross-language variant executes 3 instructions that are executed neither in the reference function nor in the same-language variant, while also avoiding 3 instructions from the original execution set. This difference in instruction subsets follows naturally from using different compiler front-ends for producing the variants’ IR code.